OpenStudy (anonymous):

Help pic below Find the scale factor of dilation If Jay(PQ) is 3'9(or 45inches)tall how tall is his dad (P'Q')

OpenStudy (anonymous):



OpenStudy (anonymous):

any idea?

OpenStudy (anonymous):

its a ratio?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so should i simplify it?

OpenStudy (anonymous):

show me wat u mean

OpenStudy (anonymous):

\[\frac{ 75 }{ 50 }=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

wat is half of 50?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

yes but the equation u just gave me said that 75 is half of fifty.

OpenStudy (anonymous):

no u would cross multiply im getting confused???

OpenStudy (anonymous):

\[\frac{ 75 }{ 50 }=\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

so that is my scale factor??

OpenStudy (anonymous):

ill be back later sorry

OpenStudy (anonymous):

ok message me when ur back

OpenStudy (anonymous):

kkk